OpenStudy (anonymous):

URGENT: a^2 + b^2 = c^2 ; solve for b

OpenStudy (anonymous):

a) subtact \(a^2\) from both sides b) take the square root

OpenStudy (anonymous):

really you should put \(\pm\) in front of the square root, but my guess is this is a triangle problem so the negative answer would not make sense

OpenStudy (anonymous):

yeah i get that you have to subtract a^2 but then what happens when you root it

OpenStudy (anonymous):

you get \[b=\sqrt{c^2-a^2}\] is all

OpenStudy (anonymous):

yeah it makes sense and i understand, but apparently it's wrong

myininaya (myininaya):

did you try the + or - option he talked about earlier ?

OpenStudy (anonymous):

would it be: \[\sqrt{c+a} \sqrt{c-a}\]

myininaya (myininaya):

I was talking about \[b=\pm \sqrt{c^2-a^2}\]

OpenStudy (anonymous):

let me try it. i got one more try to submit the answer and then it locks

myininaya (myininaya):

well is it given b is suppose to be negative or positive?

myininaya (myininaya):

or does it not say?

myininaya (myininaya):

if b>0 then b=sqrt(c^2-a^2) if b<0 then b=-sqrt(c^2-a^2)

OpenStudy (anonymous):

nope it just says "solve the equation for the indicated variable"

myininaya (myininaya):

k